A plane flies 500 miles with the wind and 350 miles against the wind in the same length of time. If the speed of the wind is 30 mph, what is the speed of the plane in still air?
Answer by mananth(16949)

Plane speed x mph
wind speed 30 mph
..
against wind x - 30 mph
with wind x + 30 mph

Distance against wind 350
Distance with wind 500
Time = same equate the time against with time with current
t=d/r
350 /(x- 30 )= 500 /(x+ 30 )
350 *(x+ 30 )= 500 *(x- 30 )
350 x + 10500 = 500 x -15000
350 x -500 x = -15000 + -10500
-150 x = -25500
/ -150
x= 170 mph ---- Plane speed
m.ananth@hotmail.ca